# Improving Alignment and Robustness with Short Circuiting

### v1

@@ Abstract
AI systems can take harmful actions and are highly vulnerable to adversarial attacks. We present an approach, inspired by recent advances in representation engineering, that “short-circuits” models as they respond with harmful outputs. Existing techniques aimed at improving alignment, such as refusal training, are often bypassed. Techniques such as adversarial training try to plug these holes by countering specific attacks. As an alternative to refusal training and adversarial training, short-circuiting directly controls the representations that are responsible for harmful outputs in the first place. Our technique can be applied to both text-only and multimodal language models to prevent the generation of harmful outputs without sacrificing utility—even in the presence of powerful unseen attacks. Notably, while adversarial robustness in standalone image recognition remains an open challenge, short-circuiting allows the larger multimodal system to reliably withstand image “hijacks” that aim to produce harmful content. Finally, we extend our approach to AI agents, demonstrating considerable reductions in the rate of harmful actions when they are under attack. Our approach represents a significant step forward in the development of reliable safeguards to harmful behavior and adversarial attacks. Code is available at github.com/blackswan-ai/short-circuiting.
@@ 1 Introduction
We propose a novel approach that fundamentally diverges from traditional defenses: instead of attempting to remove vulnerabilities to specific attacks, our approach aims to directly circumvent the ability of the model to produce the harmful output in the first place. With Short-circuiting, we make models intrinsically safer and reduce their risks by removing intrinsic model hazards—their ability to produce harmful outputs—rather than removing specific vulnerabilities with adversarial training, and rather than attempting to reduce exposure to attacks with input filters [28, 21]. Using representation engineering (RepE) [75], our method “short-circuits” the internal representations related to harmful outputs so that when a model begins to generate such an output, its internal processes are interrupted, halting completion of the generation. Because the representation used to generate a harmful output is independent of any attack capable of eliciting it, this approach is attack-agnostic, and sidesteps the need for additional training, costly adversarial fine tuning, or the use of auxiliary “guard” models. Consequently, the resulting short-circuited model can be used normally without additional computational burden, and seamlessly integrated with existing monitoring and protection mechanisms.
@@
Experimentally, we demonstrate that a short-circuiting technique, Representation Rerouting (RR), notably improves the alignment of LLMs. It enhances the harmlessness of state-of-the-art LLMs, including against against a wide array of unseen adversarial attacks, including embedding and representation-space attacks—namely, proxies for worst-case assumptions about attacker capabilities. Figure 2 and Table 1 present an overview of these results. Our method significantly outperforms standard refusal training and adversarial training, while imposing almost no penalty on standard capability. Notably, we integrate short-circuiting with additional model control methods to develop a Llama-3-8B-Instruct finetune called Cygnet. This enhanced model not only surpasses its original capabilities but also exhibits a large reduction in harmful output by approximately two orders of magnitude, even when confronted with unforeseen adversarial attacks. To the best of our knowledge, this is the first convincing demonstration of the feasibility of designing techniques that significantly advance the Pareto frontier of capability versus harmlessness for LLMs, illustrating that such trade-offs can be effectively managed. When applied to multimodal models, our results show marked increases in harmlessness. It also improves robustness against image-based attacks aimed at similarly circumventing model safeguards, again with almost no penalty on benchmarked capabilities. This remains true even in the presence of the Projected Gradient Descent (PGD) attack [37], which defenses for standalone image classifiers have been unable to achieve without a steep trade-off in accuracy. Finally, we apply short-circuiting to AI agents, illustrating its efficacy in controlling agent behaviors through evaluations on a new agent function-calling safety benchmark.
@@ 2 Related Work
Representation Engineering. As many contemporary defenses relying solely on supervising model outputs fail to achieve the desired levels of controllability and reliability, techniques that analyze and manage model’s internal representations have garnered increased attention. This includes research ranging from uncovering emergent interpretable structures in intermediate representations [75, 45, 10], to the identification and modification of embedded knowledge [47, 41, 42], as well as steering model outputs [64, 7, 32, 24, 63]. Most relevant to our work is the control vector baseline introduced in the representation engineering paper [75], which can be applied to enhance large language models’ resistance to adversarial attacks. Alongside the use of control vectors, they introduce an approach that bends representations with representation-level losses. Recent advancements extend this method to robustly unlearn hazardous knowledge [29] with a method termed RMU, demonstrating the potential of representation engineering for more complex objectives. Building on these foundations and further expanding RMU to a family of short-circuiting techniques, we design a methodology based on model representations for robust alignment and control by preventing the generation of harmful outputs.
3 Short Circuiting with Representation Engineering
@@
In this section, we introduce a novel approach aimed at mitigating the generation of harmful outputs in neural networks by inducing a new type of phenomenon called “short-circuiting.” This phenomenon can be elicited using a family of techniques designed to remap model representations related to harmful processes, redirecting them towards incoherent or refusal representations. The core objective of this method is to robustly prevent the model from producing harmful or undesirable behaviors.
@@
Our focus on generative models—such as language and multimodal agents—presents a unique opportunity. Generative models inherently involve multi-step processes through which outputs are produced. When devising an attack, adversaries must effectively exert influence across each step of the targeted processes, so each step presents an opportunity to make the model more robust to attack. This insight drives our strategy, which focuses on disrupting adversarial control of the relevant multi-step processes rather than the binary classification problem of attempting to detect the presence of an attack. Building from techniques in representation engineering [75], we accomplish this by remapping the sequence of model representations that leads to harmful outputs, directing them towards incoherent or refusal representations—namely, short-circuiting them. Moreover, by directly targeting the processes involved in generating harmful responses, our method can generalize across the diverse range of inputs that may activate those processes. Consequently, we do not need to identify all of the potential inputs that could trigger undesirable outputs, rather we only need to ensure coverage of a well defined set of such outputs.
@@
The applications of short-circuiting techniques are multifaceted. They can be utilized to prevent the generation of harmful outputs in general, as well as to prevent more narrowly tailored types of output, such as private information or copyrighted material. The approach is versatile, as it is possible to identify and remap the relevant representations in virtually any neural network architecture.
@@
The family of short-circuiting techniques is characterized by two major components: datasets and loss functions. Algorithm 1 presents a short-circuiting technique that uses Low-Rank Representation Adaptation (LoRRA) [75] which we call Representation Rerouting (RR). The remainder of this section details this approach, and how the data and chosen loss function contribute to the effectiveness of the overall method.
@@
The training data used in RR is partitioned into two sets: the Short Circuit Set and the Retain Set, each serving distinct purposes within the training process aimed at controlling harmful processes in the model. As with all representation control methods, the quality of the short-circuiting mechanism largely depends on how precisely the data can elicit the targeted representation. The Short Circuit Set is comprised of examples that yield internal representations potentially leading to harmful or undesirable behaviors, and are used to prompt the model’s short-circuiting mechanism. Conversely, the Retain Set includes examples that should not prompt short-circuiting, and are used to maintain existing desirable model representations to retain benign efficacy. While even a limited number of examples in each set can sufficiently alter the model’s behavior in a manner that generalizes beyond the training data, the resulting performance is generally improved when the training data better aligns with the domains we aim to short-circuit and retain.
For models with pre-existing refusal mechanisms, like Llama-3-Instruct, careful dataset curation is essential. Adding refusal data to the Retain Set enhances the model’s ability to correctly refuse harmful user requests and improves retention of its capabilities. Another challenge is to elicit harmful responses from models with effective refusal mechanisms. To address this, we must curate a Short Circuit set that includes text capable of bypassing the refusal mechanism and triggering harmful processes. We find that a practical approach is to remove harmful user requests while keeping the corresponding harmful assistant responses in the Short Circuit Set. These measures ensure the refusal mechanism’s integrity while allowing the model to activate its short-circuiting function correctly once the refusal is bypassed. Ablation results are detailed in Section 4.4.
@@
The accompanying losses for the datasets are the representation rerouting loss and retain loss. Denote the representation of harmful processes under the original model as and under the short-circuited model as . The rerouting loss is designed to remap representations from harmful processes to a desired target representation . Conversely, the retain loss is used to maintain representations within a retain set, which helps preserve these representations. This is often measured as the distance between the current and retain representations.
The rerouting loss can take various forms. One approach involves routing the targeted representation to a fixed random direction with a large norm, as utilized in the unlearning method RMU [29]. This is expressed as , where is a random vector and is a large constant meant to amplify the norm of the representation. However, this approach requires extensive tuning of the parameter. We also explore a variant of the random vector loss that does not necessitate hyperparameter tuning, formulated as the norm of . However, the use of a random vector is neither necessary nor optimal. Given that we want the targeted representation to be as unhelpful as possible for the harmful processes, another approach is to directly optimize the short-circuited representation to be orthogonal to the original representation responsible for harmful processes. This is given by their cosine similarity: . To avoid optimizing the similarity beyond zero, we apply a ReLU function to this objective. We find this loss to be the most intuitive and most effective in terms of achieving a balance between robustness and preserved capability. An implementation of RR using Low-Rank Representation Adaptation is shown in Algorithm 1. Additionally, one could map onto more semantically meaningful directions, such as a refusal direction or the embedding of the EOS token. We leave this to future work. Section C.1 discusses several additional design considerations.
@@ 4.1 Large Language Models
Short Circuiting. In our experimental setup, we employ similar short-circuiting and retain datasets for both the Mistral-7B-Instruct-v2 [46] and Llama-3-8B-Instruct [43] models. Detailed information on the synthetic short-circuiting set for LLMs is provided in Section A.1. The retain set for both models includes UltraChat [15], comprising instructional conversations, and XSTest [56], an exaggerated refusal dataset. Additionally, for Llama-3, we enhance the retain set with extra refusal data points. We follow the implementation of Representation Rerouting (RR) specified in Algorithm 1 and select hyperparameters based on static attack test cases from HarmBench’s validation set. More experimental details can be found in Section C.2.1.
@@
We utilize HarmBench’s LLM classifier to evaluate the attack success rate and manually verify the judgements. Detailed configurations for each attack are provided in Section C.2.2. To measure the capabilities of the short-circuited models, we evaluate our models on MTBench [73] for instruction-following abilities and on the OpenLLM Leaderboard [8] for knowledge and reasoning which includes MMLU [22], ARC-c [13], HellaSwag [70], TruthfulQA [30], Winogrande [57], and GSM8K [14]. Additionally, we follow the methodology in [5] to construct an over-refusal evaluation, described in Appendix B. For baselines, we use the original Mistral and Llama-3 Instruct models. Additionally, we include a state-of-the-art adversarially trained Mistral model, R2D2 [39], for comparison.
Results. We observe that our short-circuiting technique RR demonstrates strong generalization across a diverse range of attacks, reducing compliance rates to harmful requests by an average of with Mistral and with Llama-3. Unlike the Mistral R2D2 model, which is trained against the GCG yet shows limited generalization to various attacks, our method eliminates the need for specific attack training and focuses on hindering harmful generations. Our approach moves away from the traditional cat-and-mouse paradigm, aiming for generalization to unforeseen attacks. Additionally, the results highlight a Pareto optimal trade-off in performance. Our model exhibits high reliability against unseen attacks with a minimal compromise in capability evaluation, showing a performance dip of less than in proposed tests. This is difficult to achieve with traditional defenses. For example, the Mistral model, when adversarially trained, experiences a decline of over in the MT Bench performance. In contrast, our model leverages representation engineering principles, focusing on internal control over external supervision, enabling more targeted and fine-grained control over model behavior without adversely impacting other functionalities.
@@ 4.2 Multimodal Models
Short Circuiting. We mix the short circuit and retain datasets from Section 4.1 with a synthetic multimodal short circuit set and the retain LLaVA-Instruct set [33]. The detailed process of generating the synthetic dataset is reported in section A.2. We perform RR on LLaVA-NeXT-Mistral-7B [33]. More experimental details can be found in Section C.3.1.
Evaluation. To evaluate the robustness of multimodal short-circuited models, we generate adversarial images using a whitebox approach. Following Projected Gradient Descent [37], we perturb images with a harmful prompt to produce a target string with an affirmative assistant response. We set epsilon to and run the process for 1000 steps. As baselines, we test LLaVA-NeXT-Mistral-7B with and without a safety prompt that asks the model to avoid harmful responses. Our robustness results in Figure 3 show the percentage of harmful prompts the model complies with, labeled manually. We source a set of 133 harmful multimodal behaviors from HarmBench [39] and MM-SafetyBench [36], focusing on the most saliently harmful prompts. See Appendix C.3 for more details about the dataset’s composition. For capabilities evaluation, we follow [33] to evaluate multimodal models on LLaVA-Wild for visual chat capability and MMMU [69] for multimodal understanding capability.
Results. Figure 3 demonstrates that for multimodal models, our short-circuiting technique RR is also able to make a model significantly more robust while preserving model capabilities. Especially when subject to white-box PGD Attack, RR achieves reduction of in the compliance rate compared to the original model and compared to the safety prompt. Meanwhile, performance on MMMU and LLaVA-Wild remains within of the original model’s, as opposed to the safety prompt which causes a decrease of on LLaVA-Wild. This demonstrates that despite the ongoing challenge of achieving adversarial robustness in standalone image recognition, short-circuiting enables the larger multimodal system to reliably counter image “hijacks” [6] intended to elicit harmful outputs.

### v2

# Improving Alignment and Robustness with Circuit Breakers

@@ Abstract
AI systems can take harmful actions and are highly vulnerable to adversarial attacks. We present an approach, inspired by recent advances in representation engineering, that interrupts the models as they respond with harmful outputs with “circuit breakers.” Existing techniques aimed at improving alignment, such as refusal training, are often bypassed. Techniques such as adversarial training try to plug these holes by countering specific attacks. As an alternative to refusal training and adversarial training, circuit-breaking directly controls the representations that are responsible for harmful outputs in the first place. Our technique can be applied to both text-only and multimodal language models to prevent the generation of harmful outputs without sacrificing utility—even in the presence of powerful unseen attacks. Notably, while adversarial robustness in standalone image recognition remains an open challenge, circuit breakers allow the larger multimodal system to reliably withstand image “hijacks” that aim to produce harmful content. Finally, we extend our approach to AI agents, demonstrating considerable reductions in the rate of harmful actions when they are under attack. Our approach represents a significant step forward in the development of reliable safeguards to harmful behavior and adversarial attacks. Code is available at github.com/blackswan-ai/circuit-breakers.
@@ 1 Introduction
We propose a novel approach that fundamentally diverges from traditional defenses: instead of attempting to remove vulnerabilities to specific attacks, our approach aims to directly circumvent the ability of the model to produce the harmful output in the first place. With circuit breakers, we make models intrinsically safer and reduce their risks by removing intrinsic model hazards—their ability to produce harmful outputs—rather than removing specific vulnerabilities with adversarial training, and rather than attempting to reduce exposure to attacks with input filters [28, 21]. Using representation engineering (RepE) [76], our method connects the internal representations related to harmful outputs to circuit breakers so that when a model begins to generate such an output, its internal processes are interrupted, halting completion of the generation. Or this method is “short-circuiting” the harmful processes as one might put it. Because the representation used to generate a harmful output is independent of any attack capable of eliciting it, this approach is attack-agnostic, and sidesteps the need for additional training, costly adversarial fine tuning, or the use of auxiliary “guard” models. Consequently, the resulting model with circuit breakers can be used normally without additional computational burden, and seamlessly integrated with existing monitoring and protection mechanisms.
@@
Experimentally, we demonstrate that a circuit-breaking technique, Representation Rerouting (RR), notably improves the alignment of LLMs. It enhances the harmlessness of state-of-the-art LLMs, including against against a wide array of unseen adversarial attacks, including embedding and representation-space attacks—namely, proxies for worst-case assumptions about attacker capabilities. Figure 2 and Table 1 present an overview of these results. Our method significantly outperforms standard refusal training and adversarial training, while imposing almost no penalty on standard capability. Notably, we integrate circuit-breakering with additional model control methods to develop a Llama-3-8B-Instruct finetune called Cygnet. This enhanced model not only surpasses its original capabilities but also exhibits a large reduction in harmful output by approximately two orders of magnitude, even when confronted with unforeseen adversarial attacks. To the best of our knowledge, this is the first convincing demonstration of the feasibility of designing techniques that significantly advance the Pareto frontier of capability versus harmlessness for LLMs, illustrating that such trade-offs can be effectively managed. When applied to multimodal models, our results show marked increases in harmlessness. It also improves robustness against image-based attacks aimed at similarly circumventing model safeguards, again with almost no penalty on benchmarked capabilities. This remains true even in the presence of the Projected Gradient Descent (PGD) attack [38], which defenses for standalone image classifiers have been unable to achieve without a steep trade-off in accuracy. Finally, we apply circuit breakers to AI agents, illustrating its efficacy in controlling agent behaviors through evaluations on a new agent function-calling safety benchmark.
@@ 2 Related Work
Representation Engineering. As many contemporary defenses relying solely on supervising model outputs fail to achieve the desired levels of controllability and reliability, techniques that analyze and manage model’s internal representations have garnered increased attention. This includes research ranging from uncovering emergent interpretable structures in intermediate representations [76, 46, 10], to the identification and modification of embedded knowledge [48, 42, 43], as well as steering model outputs [65, 7, 33, 24, 64]. Most relevant to our work is the control vector baseline introduced in the representation engineering paper [76], which can be applied to enhance large language models’ resistance to adversarial attacks. Alongside the use of control vectors, they introduce an approach that bends representations with representation-level losses. Recent advancements extend this method to robustly unlearn hazardous knowledge [29] with a method termed RMU, demonstrating the potential of representation engineering for more complex objectives. Previous work has attempted to eliminate harmful circuits using bottom-up mechanistic interpretability, but these methods have proven insufficient [30]. Building on these foundations and further expanding RMU to a family of circuit-breaking techniques, we design a methodology based on model representations for robust alignment and control by preventing the generation of harmful outputs.
3 Circuit Breaking with Representation Engineering
@@
In this section, we introduce a novel approach aimed at mitigating the generation of harmful outputs in neural networks by inducing a new type of phenomenon called “circuit-breaking.” This phenomenon can be elicited using a family of techniques designed to remap model representations related to harmful processes, redirecting them towards incoherent or refusal representations. This process is reminiscent of “short-circuiting,” where harmful representations are “shorted” and intercepted by circuit breakers. The core objective of this method is to robustly prevent the model from producing harmful or undesirable behaviors.
@@
Our focus on generative models—such as language and multimodal agents—presents a unique opportunity. Generative models inherently involve multi-step processes through which outputs are produced. When devising an attack, adversaries must effectively exert influence across each step of the targeted processes, so each step presents an opportunity to make the model more robust to attack. This insight drives our strategy, which focuses on disrupting adversarial control of the relevant multi-step processes rather than the binary classification problem of attempting to detect the presence of an attack. Building from techniques in representation engineering (RepE) [76], we accomplish this by remapping the sequence of model representations that leads to harmful outputs, directing them towards incoherent or refusal representations—namely, breaking the circuit, or shorting the circuit as one might put it. Moreover, by directly targeting the processes involved in generating harmful responses, our method can generalize across the diverse range of inputs that may activate those processes. Consequently, we do not need to identify all of the potential inputs that could trigger undesirable outputs, rather we only need to ensure coverage of a well defined set of such outputs.
@@
The applications of circuit breakers are multifaceted. They can be utilized to prevent the generation of harmful outputs in general, as well as to prevent more narrowly tailored types of output, such as private information or copyrighted material. The approach is versatile, as it is possible to identify and remap the relevant representations in virtually any neural network architecture.
@@
The family of circuit-breaking techniques is characterized by two major components: datasets and loss functions. Algorithm 1 presents a circuit-breaking technique that uses Low-Rank Representation Adaptation (LoRRA) [76] which we call Representation Rerouting (RR). The remainder of this section details this approach, and how the data and chosen loss function contribute to the effectiveness of the overall method.
@@
The training data used in RR is partitioned into two sets: the Circuit Breaker Set and the Retain Set, each serving distinct purposes within the training process aimed at controlling harmful processes in the model. As with all representation control methods, the quality of the circuit breaker mechanism largely depends on how precisely the data can elicit the targeted representation. The Circuit Breaker Set is comprised of examples that yield internal representations potentially leading to harmful or undesirable behaviors, and are used to prompt the model’s circuit breaker mechanism. Conversely, the Retain Set includes examples that should not activate circuit breakers, and are used to maintain existing desirable model representations to retain benign efficacy. While even a limited number of examples in each set can sufficiently alter the model’s behavior in a manner that generalizes beyond the training data, the resulting performance is generally improved when the training data better aligns with the domains we aim to break the circuit and retain.
For models with pre-existing refusal mechanisms, like Llama-3-Instruct, careful dataset curation is essential. Adding refusal data to the Retain Set enhances the model’s ability to correctly refuse harmful user requests and improves retention of its capabilities. Another challenge is to elicit harmful responses from models with effective refusal mechanisms. To address this, we must curate a Circuit Breaker set that includes text capable of bypassing the refusal mechanism and triggering harmful processes. We find that a practical approach is to remove harmful user requests while keeping the corresponding harmful assistant responses in the Circuit Breaker Set. These measures ensure the refusal mechanism’s integrity while allowing the model to activate its circuit-breaking function correctly once the refusal is bypassed. Ablation results are detailed in Section 4.4.
@@
The accompanying losses for the datasets are the representation rerouting loss and retain loss. Denote the representation of harmful processes under the original model as and under the model with circuit breakers as . The rerouting loss is designed to remap representations from harmful processes to a desired target representation . Conversely, the retain loss is used to maintain representations within a retain set, which helps preserve these representations. This is often measured as the distance between the current and retain representations.
The rerouting loss can take various forms. One approach involves routing the targeted representation to a fixed random direction with a large norm, as utilized in the unlearning method RMU [29]. This is expressed as , where is a random vector and is a large constant meant to amplify the norm of the representation. However, this approach requires extensive tuning of the parameter. We also explore a variant of the random vector loss that does not necessitate hyperparameter tuning, formulated as the norm of . However, the use of a random vector is neither necessary nor optimal. Given that we want the targeted representation to be as unhelpful as possible for the harmful processes, another approach is to directly optimize the circuit-broken representation to be orthogonal to the original representation responsible for harmful processes. This is given by their cosine similarity: . To avoid optimizing the similarity beyond zero, we apply a ReLU function to this objective. We find this loss to be the most intuitive and most effective in terms of achieving a balance between robustness and preserved capability. An implementation of RR using Low-Rank Representation Adaptation is shown in Algorithm 1. Additionally, one could map onto more semantically meaningful directions, such as a refusal direction or the embedding of the EOS token. We leave this to future work. Section C.1 discusses several additional design considerations.
@@ 4.1 Large Language Models
Adding Circuit Breakers. In our experimental setup, we employ similar circuit breaker and retain datasets for both the Mistral-7B-Instruct-v2 [47] and Llama-3-8B-Instruct [44] models. Detailed information on the synthetic circuit breaker set for LLMs is provided in Section A.1. The retain set for both models includes UltraChat [15], comprising instructional conversations, and XSTest [57], an exaggerated refusal dataset. Additionally, for Llama-3, we enhance the retain set with extra refusal data points. We follow the implementation of Representation Rerouting (RR) specified in Algorithm 1 and select hyperparameters based on static attack test cases from HarmBench’s validation set. More experimental details can be found in Section C.2.1.
@@
We utilize HarmBench’s LLM classifier to evaluate the attack success rate and manually verify the judgements. Detailed configurations for each attack are provided in Section C.2.2. To measure the capabilities of the models with circuit breakers, we evaluate our models on MTBench [74] for instruction-following abilities and on the OpenLLM Leaderboard [8] for knowledge and reasoning which includes MMLU [22], ARC-c [13], HellaSwag [71], TruthfulQA [31], Winogrande [58], and GSM8K [14]. Additionally, we follow the methodology in [5] to construct an over-refusal evaluation, described in Appendix B. For baselines, we use the original Mistral and Llama-3 Instruct models. Additionally, we include a state-of-the-art adversarially trained Mistral model, R2D2 [40], for comparison.
Results. We observe that our circuit-breaking technique RR demonstrates strong generalization across a diverse range of attacks, reducing compliance rates to harmful requests by an average of with Mistral and with Llama-3. Unlike the Mistral R2D2 model, which is trained against the GCG yet shows limited generalization to various attacks, our method eliminates the need for specific attack training and focuses on hindering harmful generations. Our approach moves away from the traditional cat-and-mouse paradigm, aiming for generalization to unforeseen attacks. Additionally, the results highlight a Pareto optimal trade-off in performance. Our model exhibits high reliability against unseen attacks with a minimal compromise in capability evaluation, showing a performance dip of less than in proposed tests. This is difficult to achieve with traditional defenses. For example, the Mistral model, when adversarially trained, experiences a decline of over in the MT Bench performance. In contrast, our model leverages representation engineering principles, focusing on internal control over external supervision, enabling more targeted and fine-grained control over model behavior without adversely impacting other functionalities.
@@ 4.2 Multimodal Models
Adding Circuit Breakers. We mix the circuit breaker and retain datasets from Section 4.1 with a synthetic multimodal circuit breaker set and the retain LLaVA-Instruct set [34]. The detailed process of generating the synthetic dataset is reported in section A.2. We perform RR on LLaVA-NeXT-Mistral-7B [34]. More experimental details can be found in Section C.3.1.
Evaluation. To evaluate the robustness of multimodal models with circuit breakers, we generate adversarial images using a whitebox approach. Following Projected Gradient Descent [38], we perturb images with a harmful prompt to produce a target string with an affirmative assistant response. We set epsilon to and run the process for 1000 steps. As baselines, we test LLaVA-NeXT-Mistral-7B with and without a safety prompt that asks the model to avoid harmful responses. Our robustness results in Figure 3 show the percentage of harmful prompts the model complies with, labeled manually. We source a set of 133 harmful multimodal behaviors from HarmBench [40] and MM-SafetyBench [37], focusing on the most saliently harmful prompts. See Appendix C.3 for more details about the dataset’s composition. For capabilities evaluation, we follow [34] to evaluate multimodal models on LLaVA-Wild for visual chat capability and MMMU [70] for multimodal understanding capability.
Results. Figure 3 demonstrates that for multimodal models, our circuit-breaking technique RR is also able to make a model significantly more robust while preserving model capabilities. Especially when subject to white-box PGD Attack, RR achieves reduction of in the compliance rate compared to the original model and compared to the safety prompt. Meanwhile, performance on MMMU and LLaVA-Wild remains within of the original model’s, as opposed to the safety prompt which causes a decrease of on LLaVA-Wild. This demonstrates that despite the ongoing challenge of achieving adversarial robustness in standalone image recognition, circuit breakers enable the larger multimodal system to reliably counter image “hijacks” [6] intended to elicit harmful outputs.